OpenStudy (anonymous):

Which of the following is the radical expression of 4d^3/8 ?

OpenStudy (anonymous):

Your problem says which of the following, but you didn't put any "followings"

OpenStudy (anonymous):

it is in invisible mode ..............

OpenStudy (anonymous):

\[a. 4\sqrt[8]{d^3} \]\[b. 4\sqrt[3]{d^8}\] \[c.\sqrt[8]{4d^3}\] \[d. \sqrt[3]{4d^8}\]

OpenStudy (anonymous):

\[4 \times \sqrt[8]{x^3}\]

OpenStudy (anonymous):

pick a b c or d

OpenStudy (anonymous):

i cant see clearly ur options the option in which u have under root of 8 and x raised to the power of 3 is the answer

OpenStudy (anonymous):

a